OpenStudy (gabylovesyou):

Evan is painting a stool that has a cylindrical seat and four rectangular prism legs. The seat has a diameter of 14 inches and a height of 4 inches. Each leg is 4 inches by 4 inches by 18 inches. How much area will Evan have to paint? Use 3.14 for TT. Show your work.

jimthompson5910 (jim_thompson5910):

Do you have an answer that you need to check or do you need me to go through the steps?

OpenStudy (gabylovesyou):

Can u go through the steps! I didn't know I missed 2 questions and i don't get them :S

jimthompson5910 (jim_thompson5910):

Area of the top of the stool: A = pi*r^2 A = (3.14)*(7)^2 ... The diameter is 14 in, so the radius is 14/2 = 7 in A = 153.86 square inches ----------------------------------- Lateral surface area of the top of the stool A = 2pi*r*h A = 2*(3.14)*7*4 A = 175.84 square inches ------------------------------------- Surface area of one rectangular prism leg (ignoring the top face) SA = 2*(LW+WH+LH) - LW SA = 2*(4*4+4*18+4*18) - 4*4 SA = 2*(4*4+4*18+4*18) - 4*4 SA = 304 There are 4 of these legs. So the total surface area of the legs is 4*304 = 1216 square inches. -------------------------------------- Now add up the area of the top, the sides of the stool, and the legs of the stool to get 153.86 + 175.84 + 304 = 633.7 So Evan will have to paint 633.7 square inches total.

OpenStudy (anonymous):

you need more help gaby?

OpenStudy (gabylovesyou):

Hi @dpaInc i might need more help! if i do i will let u know :D

OpenStudy (anonymous):

ok...jim did a great job btw...

jimthompson5910 (jim_thompson5910):

thx dpaInc

OpenStudy (gabylovesyou):

yes he did :D

jimthompson5910 (jim_thompson5910):

lol glad I could help you out

OpenStudy (gabylovesyou):

i feel so embarrassed u have been helping me out alot! sorry

jimthompson5910 (jim_thompson5910):

it's no trouble, as long as I'm truly helping you I don't mind

OpenStudy (gabylovesyou):

ohh ok and yes u are! DEFINITELY helping :D

jimthompson5910 (jim_thompson5910):

lol great

OpenStudy (anonymous):

look what i found lmao @Gabylovesyou